OpenStudy (anonymous):

Directions: Find the value of n for each regular n-gon described. Question: Each interior angle of the regular n-gon has a measure of 156 degrees. Please show work!!! thank you:)

OpenStudy (anonymous):

So are you aware that the formula of interior angle of n-gon => \[\frac{ (n-2)180 }{ n }\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

if i'm trying to find n? then the equation is (180n-360)/n ?

OpenStudy (anonymous):

how does this equation help me find n?

OpenStudy (anonymous):

Yes. Now you said each interior angle = 156. So - \[\frac{ 180n-360 }{ n } = 156\]

OpenStudy (anonymous):

180n - 360 = 156n. 180n - 156n = 360. 24 n = 360.

OpenStudy (anonymous):

thank you!